Solve the following equations. 
16 = y/15
Solve the following equations. 
16 = y/15

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

1.6 = y/1.5
y/1.5 = 1.6
y = 1.6 × 1.5
y=2.4

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve rhis question as follows :
16 = Y/15
Y = 16 × 15
Y = 240
 
Hope it helps!
Thank you ✌️